OpenStudy (anonymous):

What is the value of the expression 7^-3 A. -21 B. -1/343 C. 1/343 D. 343

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \bf a^{-4}= \frac{1}{a^{4} } }\) \(\Huge\color{blue}{ \bf 7^{-3}= ? }\) you tell me

OpenStudy (anonymous):

i dont understand how to do any of this....

OpenStudy (solomonzelman):

you are just saying that if a number is raise to negative exponent, that it is the same as 1 divided by this number raised to a Positive exponent. For example \(\Huge\color{blue}{ \bf 3^{-5}= \frac{1}{3^{5} }=\frac{1}{125 } }\)

OpenStudy (anonymous):

so........ im gonna say its C..

OpenStudy (solomonzelman):

sorry not 125... 243

OpenStudy (solomonzelman):

yes, you are correct.

OpenStudy (anonymous):

oh okay. thank you